# On the Approximation Power of SiLU Networks: Exponential Rates and Depth Efficiency

### v1

@@ 3.1 SiLU is robust for polynomial approximation
For every $B>0,m\geq 3,$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{m}>0$, $\omega_{m}>1$ and a family of feedforward SiLU neural networks $\{P_{m,k}\}_{k\in\mathbb{N}}$, with at most $2m$ hidden layers with $4$ neurons each, such that for $k\geq\max\left\{\left\lceil\frac{\ln C_{m}-\ln\epsilon}{2\ln\omega_{m}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{m}}\right\rceil\right\}$
$\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq\epsilon$
@@
This theorem is proven by induction on the degree $m$, building upon the multiplier network $M_{k}$ from the previous theorem. The base cases for $m=0,1,2$ are constructed directly (as a constant, the identity function, and the squaring network $Q_{k}$, respectively). The inductive step approximates $x^{m+1}$ by using the network $M_{k}$ to multiply the approximation of $x^{m}$ by $x$.
The resulting network $P_{m,k}$ for $m\geq 2$ has a recursive structure. It can be expressed as a composition of networks where each stage $i$ computes a product, leading to the final explicit architecture:
$P_{m,k}(x)=\gamma_{m}\cdot\text{SiLU}\left(W_{m}z_{m}+b_{m}\right)$ (3)
with the parameters defined as:
$W_{i}=\beta^{k}\begin{bmatrix}+1&+1\\ -1&-1\\ +1&-1\\ -1&+1\end{bmatrix},\quad\gamma_{i}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1\\ +1\\ -1\\ -1\end{bmatrix}^{\top},\quad b_{i}=[a,a,a,a]^{\top}.$ (4)
The input vector is $z_{i}=[x,\gamma_{i-1}\cdot\text{SiLU}(Wz_{i-1}+b_{i-1})]^{\top},\ \forall\ i\in\{2,3,\cdots,m\}$ such that $z_{1}=[x,1]^{\top}.$ This architecture ensures the network maintains a depth of $\mathcal{O}(m)$ and a size of $\mathcal{O}(1)$. The detailed proof, can be found in Appendix A.3
Figure 6: Approximation of the function $x^{7}$ by the functions $P_{m,k}$ (Theorem 3).
@@ 3.1.1 A Natural Recurrent Neural Network (RNN) Construction
$\mathbf{W}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ +1\\ -1\end{bmatrix},\ \mathbf{U}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ -1\\ +1\end{bmatrix},\ \mathbf{b}_{h}=[a,a,a,a]^{\top},\ \mathbf{W}_{y}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}$ (6)
And the initial states: $y_{0}=1$ (We can consider that the first operation $z_{1}=[x,1]^{\top}$ amounts to initializing the hidden state $h_{1}$ with $y_{0}=1$).
Figure 7: Approximation of the $x^{7}$ function by a recurrent neural network (5) with parameters given by (6)($k=3$).
$P_{m,k}$ is indeed a shallow RNN (a $2m$ hidden layer of $4$ neurons per step) whose parameters are fixed, shared, and not learned in this context, as they were chosen analytically to compute a specific function ($x^{m}$) rather than learned from data. ∎
This final result demonstrates that the ability to approximate efficiently monomials functions, as established in the previous theorem, directly extends to the approximation of arbitrary polynomials. The following theorem shows that a SiLU network can approximate any polynomial with exponential accuracy, where the complexity of the network scales linearly with the degree of the polynomial.
@@
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be a polynomial of degree $M$. For every $B>0$ and target accuracy $\epsilon\in(0,1)$, there exist constants $C_{M}>0$, $\omega_{M}>1$ and a family of SiLU neural network $\{\mathcal{N}_{M,k}\}_{k\in\mathbb{N}}$, with at most $2M$ hidden layers with $4$ neurons each, such that for $k=\left\{\left\lceil\frac{\ln(C_{M}\cdot\max_{m}|a_{m}|)-\ln\epsilon}{2\ln\omega_{M}}\right\rceil,\left\lceil\frac{\ln 2B}{2\ln\omega_{M}}\right\rceil\right\}$
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\epsilon$
@@
This result is a direct consequence of the previous theorem on monomial approximation. The network $\mathcal{N}_{M,k}$ is constructed as a linear combination of the approximating networks $P_{m,k}$ for each monomial $x^{m}$, using the coefficients $a_{m}$ of the polynomial. The exponential error bound follows from the triangle inequality and the individual error bounds of each $P_{m,k}$. The depth of the network is dominated by the deepest monomial network, which is $O(M)$, while the size is $O(1)$ due to parameter sharing and the fact that the monomial networks $P_{M,k}$ that approximates monomial $x^{M}$ computes recursively approximated monomial of real monomial functions $x^{m},m\leq M.$ The complete proof, can be found in Appendix A.4.
Figure 8: Approximation of the function $x^{2}+x+1$ by the functions $\mathcal{N}_{M,k}$ (Theorem 4).
@@
$\frac{1}{n}\|\mathbf{\hat{y}}-\mathbf{P}\mathbf{c}\|_{2}.$
Gradient descent can be applied to learn $\mathbf{c}$. If the matrix $\mathbf{P}^{\top}\mathbf{P}$ is invertible, the exact solution can be obtained in one step:
$\mathbf{c^{*}}=(\mathbf{P}^{\top}\mathbf{P})^{-1}\mathbf{P}^{\top}\mathbf{y}.$
Corollary 2 (Universal Polynomial Approximation via SiLU Recurrent Networks).
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be any polynomial of degree $M$. There exists a parameter-sharing recurrent neural network with SiLU activation that computes:
$\mathcal{N}_{M,k}(x)=\sum_{m=0}^{M}c_{m}y_{m}$
where the hidden states $y_{m}$ are computed sequentially through the recurrence:
$y_{0}$ $=1$
$\mathbf{h}_{m}$ $=\mathrm{SiLU}\left(\mathbf{W}_{h}x+\mathbf{U}_{h}y_{m-1}+\mathbf{b}_{h}\right)$
$y_{m}$ $=\mathbf{W}_{y}\mathbf{h}_{m}\quad\text{for }m=1,2,\ldots,M$
with fixed shared parameters:
$\mathbf{W}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ +1\\ -1\end{bmatrix},\ \mathbf{U}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ -1\\ +1\end{bmatrix},\ \mathbf{b}_{h}=\begin{bmatrix}a\\ a\\ a\\ a\end{bmatrix},\ \mathbf{W}_{y}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1&+1&-1&-1\end{bmatrix}$
such that
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\epsilon.$
Least‑squares realisation. In the sense of the $\|\cdot\|_{2}$ norm, the optimal coefficients $\mathbf{c}^{\ast}=(c_{0}^{\ast},\ldots,c_{M}^{\ast})$ for approximating $N(x)$ from data $\{(x_{j},y_{j})\}_{j=1}^{n}$ with $y_{j}=N(x_{j})$ can be obtained in closed form. Provided the matrix $\mathbf{P}^{\top}\mathbf{P}$ is invertible, the solution is given by:
$\mathbf{c}^{\ast}=(\mathbf{P}^{\top}\mathbf{P})^{-1}\mathbf{P}^{\top}\mathbf{y},$
where the design matrix $\mathbf{P}\in\mathbb{R}^{n\times(M+1)}$ has entries $\mathbf{P}_{j,m}=y_{m}(x_{j})$. This approximation satisfies the error bound:
$\|\mathcal{N}_{M,k}(x)-N(x)\|_{2}\leq\epsilon.$
Proof.
The result follows directly from Theorem 4 by observing that the recurrent network’s sequential computation naturally generates the monomial basis $\{1,x,x^{2},\ldots,x^{M}\}$ through its hidden states $y_{m}$ (Figure 7). The closed-form solution for the coefficients is obtained by recognizing that the approximation problem reduces to linear regression in the feature space spanned by the network’s hidden states. The error bound is inherited from the monomial approximation bounds of the constituent networks. ∎
Remark 2.
This corollary reveals a remarkable property: a single recurrent network with fixed parameters can serve as a universal polynomial feature extractor. The network’s architecture effectively implements a polynomial basis expansion, where:
  • The depth ($M$ steps) determines the maximum polynomial degree
  • The hidden dimension (4 units) determines the approximation quality
  • The coefficients $\mathbf{c}$ can be learned efficiently via convex optimization
This provides a computationally efficient alternative to traditional polynomial regression while maintaining interpretability through the polynomial coefficients.
Building upon the previous theorem 3 that demonstrated the approximation of the monomial $x^{m}$ using a deep architecture with $m$ hidden layers of four neurons each, the current theorem represents a significant improvement by reducing the architecture to a single hidden layer of $m+1$ neurons. This structural optimization is achieved by systematically leveraging the compositional properties and finite difference calculus of the SiLU activation function, which allows the polynomial approximation to be condensed into a single non-linear transformation while maintaining exponential convergence. The shift from depth to a moderate increase in width offers a practical advantage for implementations where depth-related constraints, such as vanishing gradients or computational complexity, are more critical than the width of individual layers.
3.2 A single hidden layer SiLU Network can approximate $x^{m}$ with arbitrary precision on any compact set ?
Theorem 5.
For any $B>0$ and target accuracy $\epsilon\in(0,1)$ , there exist constants $C_{1}>0$, $\omega_{1}>1$ and a family of feedforward SiLU neural networks $\{Q^{m}_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $m+1$ neurons each, such that for $k=\max\left\{\left\lceil\frac{\ln mB-\ln 2}{\ln\omega}\right\rceil,\left\lceil\frac{ln\epsilon-\ln C_{1,m}}{\ln\omega}\right\rceil\right\}$
$\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq\epsilon,$
Proof.
The complete proof is given in Appendix A.5. The construction begins by defining a linear combination of SiLU activations, $g_{a,m}(x)$, whose Taylor expansion is analyzed using finite difference calculus (Appendix B). This reveals that $g_{a,m}(x)=K_{m}(a)x^{m}+\text{(higher-order terms)}$. The function $f_{a,m}(x)=g_{a,m}(x)/K_{m}(a)$ thus provides an $m$-th order approximation to $x^{m}$. An iterative refinement is then applied: defining $Q^{m}_{k}(x)=\beta^{-m}Q^{m}_{k-1}(\beta x)$ with $Q^{m}_{0}=f_{a,m}$. It is shown that this process exponentially suppresses the higher-order error terms. A detailed analysis of the Taylor coefficients and the application of Cauchy’s integral formula (Appendix C) yield an explicit bound on the approximation error, $\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq C_{1,m}\omega^{-k}$, proving that a target accuracy $\epsilon$ is achieved for a suitable parameter $k$. The network $Q^{m}_{k}$ is explicitly parameterized, confirming it is a shallow network with $\mathcal{O}(m)$ parameters. Then we have
$Q^{m}_{k}(x)=\gamma_{m}\cdot\text{SiLU}(W_{m}x+b_{m})$ (7)
with parameters such that the weights
$W=\beta^{k}\,[w_{0},w_{1},\dots,w_{m}]^{\top},\gamma=\frac{\beta^{-mk}}{K_{m}(a)}[c_{0},c_{1},\dots,c_{m}],$
and the bias
$b=a\mathbf{1}_{m+1}.$
with
$w_{j}=j-\frac{m}{2},\quad c_{j}=(-1)^{m-j}\binom{m}{j}$,
and where $\mathbf{1}_{m+1}$ is the column vector of size $m+1$ containing ones. ∎
Remark 3.
The condition $K_{m}(a)\neq 0$ draws our attention when $a=0.$ Indeed in this case,
$K_{m}(0)=$ $0\sigma^{(m)}(0)+m\sigma^{(m-1)}(0)$
$=m\sigma^{(m-1)}(0)$
$=mf^{(m-1)}(0),\quad\text{with}\quad f(x)=\sigma(x)-1/2$
Noting that f is an odd function, it follows that $f^{(m-1)}(-x)=(-1)^{m}f^{(m-1)}(x).$ Thus, if $m$ is odd and $m\geq 3$
$f^{(m-1)}(0)=(-1)f^{(m-1)}(0)\implies f^{(m-1)}(0)=\sigma^{(m-1)}(0)=K_{m}(0)=0.$
We have shown that a SiLU network can approximate uniformly on a compact set any monomial $x^{m}$, for an arbitrary degree $m$, with controlled depth and width. Since monomials form a natural generating family for spaces of regular functions through polynomial expansions, this polynomial-approximation capability will serve as the foundation for the analysis that follows. In the next section, we extend this elementary result to the fast approximation of continuous functions, and then to functions with higher regularity, by SiLU neural networks.
3.3 Fast approximation of continuous functions and general smooth functions by SiLU Neural Networks
We establishes the approximation capability of SiLU networks for continuous functions, quantifying the efficiency of this approximation. The following result demonstrates that any continuous function on a compact interval can be approximated with arbitrary precision $\varepsilon$ by a SiLU network, and it provides explicit bounds on the required depth and size of the network to achieve this error. The key insight is the combination of classical functions (like polynomial) approximation (Weierstrass theorem) with the efficient neural approximation of polynomials developed in the previous theorems. In this section, we approximate functions on a compact set using two approaches: the polynomial approach and the step function approach.

### v2

@@ 3.3.1 Deep Network Construction with Constant Width
For every $B>0,m\geq 2,$ and target accuracy $\varepsilon\in(0,1)$, there exist constants $C_{m}>0$, $\omega_{m}>1$ and a family of feedforward SiLU neural networks $\{P_{m,k}\}_{k\in\mathbb{N}}$, with at most $2m$ hidden layers with $4$ neurons each, such that for $k\geq\left\lceil\frac{\ln C_{m}-\ln\varepsilon}{2\ln\omega_{m}}\right\rceil$
$\sup_{|x|\leq B}|P_{m,k}(x)-x^{m}|\leq\varepsilon$
@@
By induction on $m$. For $m=2$ use Theorem 1 directly.
Assume $P_{m,k}$ satisfies $|P_{m,k}(x)-x^{m}|\leq C_{m}\omega_{m}^{-2k}$ for $|x|\leq B$ with $\mathcal{O}(m)$ depth, $\mathcal{O}(1)$ size. Define
$P_{m+1,k}(x)=M_{k}(P_{m,k}(x),x),$
where $M_{k}$ is the product network from Theorem 2. Since both inputs are bounded by $B^{\prime}=B+\max_{|x|\leq B}|P_{m,k}(x)|$, we have
$|P_{m+1,k}(x)-x^{m+1}|$
$\leq|M_{k}(P_{m,k}(x),x)-P_{m,k}(x)x|+|P_{m,k}(x)-x^{m}||x|$
$\leq(C_{2}+B^{\prime}C_{m})\omega_{m+1}^{-2k},$
with $\omega_{m+1}=\min\{\omega_{2},\omega_{m}\}$. Depth increases by $\mathcal{O}(1)$, size remains $\mathcal{O}(1)$. The choice
$k\geq\left\lceil\frac{\ln C_{m+1}-\ln\varepsilon}{2\ln\omega_{m+1}}\right\rceil$
yields error $\varepsilon$. ∎
In practice (Fig. 6), we achieve an approximation of the function $x^{7}$ using the NN $P_{m,k}$ with parameters $a=0$ and $\beta=0.27$.
Figure 6: Approximation of the function $x^{7}$ by the functions $P_{m,k}$ with $a=0$ and $\beta=0.27$ (Theorem 3).
3.3.2 Shallow Construction for Arbitrary Monomials
While Theorem 3 approximates $x^{m}$ using a deep network with $\mathcal{O}(m)$ layers, we now present a more efficient shallow construction requiring only a single hidden layer. This architectural improvement is made possible by exploiting the finite difference properties of the SiLU activation function, which allows the polynomial approximation to be compressed into a single nonlinear transformation while maintaining exponential convergence. The shift from depth to moderate width offers practical advantages where depth-related issues like vanishing gradients are a concern.
Theorem 4.
For any $B>0$ and target accuracy $\varepsilon\in(0,1)$ , there exist constants $C_{1,m}>0$, $\omega>1$ and a family of feedforward SiLU neural networks $\{Q^{m}_{k}\}_{k\in\mathbb{N}}$, with one hidden layer with $m+1$ neurons each, such that for $k=\left\lceil\frac{\ln C_{1,m}-ln\varepsilon}{\ln\omega}\right\rceil$
$\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq\varepsilon,$
Proof.
For $m\geq 1$ and $a\in\mathbb{R}$, define
$g_{a,m}(x)=\sum_{j=0}^{m}(-1)^{m-j}\binom{m}{j}\mathrm{SiLU}\!\bigl((j-\tfrac{m}{2})x+a\bigr),$
where $\mathrm{SiLU}(t)=t\sigma(t)$ and $\sigma(t)=(1+e^{-t})^{-1}$.
Expanding $\sigma$ in Taylor series around $a$ gives
$\mathrm{SiLU}\!\bigl((j-\tfrac{m}{2})x+a\bigr)$ $=a\sigma(a)+\sum_{n\geq 1}\frac{(j-\tfrac{m}{2})^{n}}{n!}\bigl[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\bigr]x^{n}.$
Summing over $j$ with alternating binomial coefficients cancels all terms with $n<m$ (the discrete $m$-th difference of a polynomial of degree <m<m vanishes). Hence
$g_{a,m}(x)=K_{m}(a)x^{m}+\sum_{n>m}A_{n}x^{n},$
where $K_{m}(a)=a\sigma^{(m)}(a)+m\sigma^{(m-1)}(a)\neq 0$ and $A_{n}=\frac{\Delta^{m}p_{n}(0)}{n!K_{m}(a)}\bigl[a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)\bigr]$ with $p_{n}(t)=(t-\tfrac{m}{2})^{n}$.
Set $f_{a,m}(x)=g_{a,m}(x)/K_{m}(a)=x^{m}+\sum_{n>m}A_{n}x^{n}$. For $\beta\in(0,1)$ define
$Q^{m}_{k}(x)=\beta^{-mk}f_{a,m}(\beta^{k}x)=x^{m}+\sum_{n>m}A_{n}\beta^{k(n-m)}x^{n}.$
Using the bound $|\Delta^{m}p_{n}(0)|\leq 2^{m}(m/2)^{n}$ (B) and $|a\sigma^{(n)}(a)+n\sigma^{(n-1)}(a)|\leq C_{a}(2/\pi)^{n}n!$ (A), we obtain $|A_{n}|\leq C_{1}(m/\pi)^{n}$ with $C_{1}=2^{m}C_{a}/|K_{m}(a)|$.
Hence for $|x|\leq B$,
$|Q^{m}_{k}(x)-x^{m}|$ $\leq C_{1}\beta^{-mk}\sum_{n>m}\bigl(\tfrac{mB}{\pi}\beta^{k}\bigr)^{\!n}$
$\leq C_{1}\beta^{-mk}\bigl(\tfrac{mB}{\pi}\beta^{k}\bigr)^{m+1}\sum_{n\geq 0}\bigl(\tfrac{mB}{\pi}\beta^{k}\bigr)^{\!n}$
$\leq C_{1}\beta^{k}\bigl(\tfrac{mB}{\pi}\bigr)^{m+1}\frac{1}{1-\frac{mB\beta^{k}}{\pi}}$
$\leq C_{1,m}\,\beta^{k},$
where $C_{1,m}=2C_{1}\left(\frac{mB}{\pi}\right)^{m+1}$ provided $\beta^{k}mB/\pi<1/2$. Thus
$\sup_{|x|\leq B}|Q^{m}_{k}(x)-x^{m}|\leq C_{1,m}\omega^{-k},$
with $\omega=\beta^{-1}$. Choosing $k=\mathcal{O}(\ln 1/\varepsilon)$ yields error $\varepsilon$. The network $Q^{m}_{k}$ is a single hidden layer SiLU network with $m+1$ neurons:
$Q^{m}_{k}(x)=\gamma_{m}\cdot\mathrm{SiLU}(W_{m}x+b_{m}),$
where $W_{m}=\beta^{k}[w_{0},\dots,w_{m}]^{\top}$ with $w_{j}=j-\tfrac{m}{2}$, $\gamma_{m}=\beta^{-mk}K_{m}(a)^{-1}[c_{0},\dots,c_{m}]$ with $c_{j}=(-1)^{m-j}\binom{m}{j}$, and $b_{m}=a\mathbf{1}_{m+1}$, where $\mathbf{1}_{m+1}=[1,1,\dots,1]^{\top}\in\mathbb{R}^{m+1}$ is the all-ones vector of dimension $m+1$. ∎
Remark 1.
For $a=0$, we have $K_{m}(0)=mf^{(m-1)}(0)$ where $f(x)=\sigma(x)-1/2$ is odd. Consequently, if $m$ is odd and $m\geq 3$, then $f^{(m-1)}(0)=0$ and thus $K_{m}(0)=0$.
@@ 3.3.3 A Natural Recurrent Neural Network (RNN) Construction
$\mathbf{W}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ +1\\ -1\end{bmatrix},\ \mathbf{U}_{h}=\beta^{k}\begin{bmatrix}+1\\ -1\\ -1\\ +1\end{bmatrix}\ \mathbf{W}_{y}=\frac{\beta^{-2k}}{4K(a)}\begin{bmatrix}+1\\ +1\\ -1\\ -1\end{bmatrix}$, (4)
and $\mathbf{b}_{h}=\begin{bmatrix}a&a&a&a\end{bmatrix}$.
The initial state is $y_{0}=1$. $P_{m,k}$ is indeed a shallow RNN (a $2m$ hidden layer of $4$ neurons per step) whose parameters are fixed, shared, and not learned in this context, as they were chosen analytically to compute a specific function ($x^{m}$) rather than learned from data. ∎
We implement in practice an approximation of $x^{7}$ with a recurrent neural network Eq. (3) (parameters: $k=3$, $a=0$, $\beta=0.27$ as per Eq. (4)). The Figure 7 reveals that each hidden layer approximates a monomial of increasing degree.
Figure 7: Approximation of the $x^{7}$ function by a recurrent neural network (3) with parameters given by Eq. (4)($k=3$, $a=0$ and $\beta=0.27$).
3.4 Approximation of Arbitrary Polynomials
The efficient approximation of monomials established in the previous theorems directly extends to arbitrary polynomials. The following result shows that any polynomial can be approximated by a SiLU network with exponential accuracy, where the network depth scales linearly with the polynomial degree while the width remains constant.
@@ 3.4 Approximation of Arbitrary Polynomials
Let $N(x)=\sum_{m=0}^{M}a_{m}x^{m}$ be a polynomial of degree $M$. For every $B>0$ and target accuracy $\varepsilon\in(0,1)$, there exist constants $C_{M}>0$, $\omega_{M}>1$ and a family of SiLU neural network $\{\mathcal{N}_{M,k}\}_{k\in\mathbb{N}}$, with at most $2M$ hidden layers with $4$ neurons each, such that for $k=\left\lceil\frac{\ln(C^{\prime}_{M}\cdot\max_{m}|a_{m}|)-\ln\varepsilon}{2\ln\omega^{\prime}_{M}}\right\rceil$
$\sup_{|x|\leq B}|\mathcal{N}_{M,k}(x)-N(x)|\leq\varepsilon$
@@
For each $m=0,\dots,M$, let $P_{m,k}$ be the SiLU network approximating $x^{m}$ on [B,B][-B,B] with error $\leq C_{m}\omega_{m}^{-2k}$, depth $\mathcal{O}(m)$, and constant size (Theorem 3).
Define $\mathcal{N}_{M,k}(x)=\sum_{m=0}^{M}a_{m}P_{m,k}(x)$. Then
$|\mathcal{N}_{M,k}(x)-N(x)|$ $\leq\sum_{m=0}^{M}|a_{m}|\,|P_{m,k}(x)-x^{m}|$
$\leq\Bigl(\max_{m}|a_{m}|\Bigr)\sum_{m=0}^{M}C_{m}\omega_{m}^{-2k}.$
Setting $C^{\prime}_{M}=(\max_{m}|a_{m}|)\sum_{m=0}^{M}C_{m}$ and $\omega^{\prime}_{M}=\min_{m}\omega_{m}$ yields
$|\mathcal{N}_{M,k}(x)-N(x)|\leq C^{\prime}_{M}\omega_{M}^{{}^{\prime}-2k}.$
The networks $P_{m,k}$ can be evaluated jointly in a single recurrent structure of depth $\mathcal{O}(M)$ and constant width; adding the final linear combination preserves these bounds. Hence $\mathcal{N}_{M,k}$ has depth $\mathcal{O}(M)$, constant size, and achieves error $\varepsilon$ for
$k=\left\lceil\frac{\ln(C^{\prime}_{M}\cdot\max_{m}|a_{m}|)-\ln\varepsilon}{2\ln\omega^{\prime}_{M}}\right\rceil$
In the same spirit, to approximate $x^{2}+x+1$ using the neural network $\mathcal{N}_{2,k}$ with parameters $k=3$, $a=0$, $\beta=0.27$, We employ the network $P_{2,k}$ to approximate the monomial $x^{2}$. According to Corollary 1, since $P_{2,k}$ is an RNN, its hidden and output layers approximate the constant function $1$, the identity function $x$, and $x^{2}$ respectively. Consequently, the sum of the outputs of these layers accurately approximates the polynomial $x^{2}+x+1$ (Fig 8).
Figure 8: Approximation of the function $x^{2}+x+1$ by the NN $\mathcal{N}_{M,k}$ (Corollary 1).
@@
$\|\mathbf{\hat{y}}-\mathbf{P}\mathbf{c}\|_{2}.$
The efficient approximation of monomials established above serves as a fundamental building block for approximating more general function classes. We now extend these results to continuous functions and then to functions with higher regularity, deriving corresponding approximation rates for SiLU networks.
3.5 Approximation of Continuous Functions
This subsection establishes approximation rates for arbitrary continuous functions using SiLU networks. Building on the efficient polynomial approximation results from Section 3, we present two methods: (i)a step-function approach whose complexity is governed by the modulus of continuity of the target function, and (ii)direct polynomial approximation via the Weierstrass theorem.
The following theorem formalizes the step-function approach, which provides explicit depth and size bounds in terms of the desired accuracy $\varepsilon$ and the function’s regularity.